OpenStudy (anonymous):

Suppose an investment of $2,300 doubles in value every decade. The function gives the value of the investment after x decades. f(x)=2300*2^x How much is the investment worth after 2 decades? (1 point)$2,355,200 $92,000 $46,000 $9,200

OpenStudy (anonymous):

Can one of you help me?(:

OpenStudy (saifoo.khan):

Replace x with 2.

OpenStudy (whpalmer4):

f(x) = 2300*2^x x = number of decades = 2 f(2) = 2300*2^2 = 2300*4 = ?

OpenStudy (anonymous):

I ried that, but I didnt know if it was right or not. Bc I got the first one

OpenStudy (anonymous):

Why would I multiply by 4 at the end?

OpenStudy (whpalmer4):

2^2 = (2)(2) = 4 so $2300*4 =

OpenStudy (anonymous):

9,200?

OpenStudy (whpalmer4):

Do you understand what 2^x means?

OpenStudy (whpalmer4):

or \[2^x\]

OpenStudy (anonymous):

Yeah

OpenStudy (whpalmer4):

Okay, that's why you multiply $2300 by 4 :-) What if the number of decades was 5, what would the answer be?

OpenStudy (anonymous):

Ughhh, i am so confused. I thought I was supposed to do it like this. 2300*2^2 and then get 9200! Right?

OpenStudy (whpalmer4):

No, I was asking for the case where the number of decades is 5. x=5

OpenStudy (whpalmer4):

So the formula would be f(5) = 2300*2^5 = ?

OpenStudy (anonymous):

I know, but I am talking about this problem. Was I right abpout what I just said?

OpenStudy (whpalmer4):

Yes, for this problem the answer is 9200. I was attempting to make sure that you understood what we actually did, that's why I asked you for the answer if the number of decades is 5.

OpenStudy (anonymous):

did you get it right @Heyy_Girl_Heyy ???